OpenStudy (vane11):

Need Help Please! I know what I'm supposed to do but keep getting the wrong answer =/ Use the laws of logarithms to expand and simplify the expression. ln((x^2)/(square root x (1+4)^4))

OpenStudy (vane11):

\[\ln ((x ^{2})/(\sqrt{x} (1+x)^{4})\] I messed up the first time I wrote the equation this is the correct one

ganeshie8 (ganeshie8):

\[\large \ln \left(\dfrac{x^2}{\sqrt{x}(1+x)^4}\right)\]

ganeshie8 (ganeshie8):

like that ?

OpenStudy (vane11):

Yeah I tried making it look nicer haha

ganeshie8 (ganeshie8):

good, do u knw any log rules ?

OpenStudy (vishweshshrimali5):

@Vane11, Ganeshi means these: (1)\(\large{\ln(\cfrac{a}{b}) = \ln{a} - \ln{b}}\) (2) \(\large{\ln(ab) = \ln {a} + \ln{b}}\) (3) \(\large{\ln(a^b) = b\ln a}\)

OpenStudy (vane11):

Yes I do :)

OpenStudy (vishweshshrimali5):

Good: then apply them in the given order.

OpenStudy (vane11):

the website is acting strange takes forever and freezes sorry for the delays btw

OpenStudy (vishweshshrimali5):

Yeah. It happens even with me. :)

OpenStudy (vishweshshrimali5):

So, first use the law 1 to change the division into subtraction of 2 logarithms.

OpenStudy (vishweshshrimali5):

For example: See: \[\large \ln \left(\dfrac{x^2}{\sqrt{x}(1+x)^4}\right)\] \[\large{ = \ln (x^2) - \ln(\sqrt{x}(1+x)^4)}\]

OpenStudy (vane11):

so then 2lnx for the first part right?

OpenStudy (vishweshshrimali5):

yes

OpenStudy (vane11):

I know sqrt is exponent 1/2 but idk if im supposed to separate the second half into two sections: -1/2lnx - 4ln(1+x) I feel that's wrong though

OpenStudy (vishweshshrimali5):

well I don't think there is anything wrong in it

OpenStudy (vane11):

Yeah I put that answer down and its marked wrong, I've tried playing with the parenthesis and leaving the second half as is as well as the first but no combination has worked for me

OpenStudy (vishweshshrimali5):

\[\large{2\ln{x} - \cfrac{1}{2}\ln{x} - 4\ln(1+x)}\] \[\large{= \cfrac{3}{2}\ln{x} - 4\ln(1+x)}\]

OpenStudy (vane11):

hm good point like terms hadn't thought of that

OpenStudy (vishweshshrimali5):

try this

OpenStudy (vane11):

it worked! I guess it was the expand AND simplify that I was leaving out, I'd expand it but forgot about combining like terms, thanks! definitely a good lesson to remember for the exam :)

OpenStudy (vishweshshrimali5):

:)